The sum of the ages of Sandy and Amy is 55. Amy is 9 years younger than Sandy. How old is Sandy?
Answer by ptth176(50)

the age of Sandy = x
the ages of Amy = y
The sum of the ages of Sandy and Amy is 55=>x+y=55(1)
Amy is 9 years younger than Sandy=>x=y+9(2)
(1)(2)=>y+y+9=55
=>2y=46
=>y=23
=>x=32
=>Sandy is 32 years old